Suppose Alice and Drake are doing Physics and Math homework. In one hour, Alice can solve 5 Physics questions or 10 Math questions, or any other combination lying on the line between those two points. In the same amount of time, Drake can solve 10 Physics questions or 15 Math questions, or any other combination lying on the line between those two points.

a.

What is the opportunity cost of solving 1 Math question for Alice? What about for Drake?

b.

What is the opportunity cost of solving 1 Physics question for Alice? What about for Drake?

c.

Who has the absolute advantage in doing Physics homework and Math homework?

d.

Suppose they work together. Who will specialize in Math problems and who will specialize in Physics problems?

e.

If they were to trade with each other, what is the price range (in terms of number of Math problems) acceptable to both Alice and Drake for a Physics question.


1
Expert's answer
2020-10-07T08:53:07-0400

a. The opportunity cost of solving 1 Math question for Alice is 5/10 = 0.5 Physics question. The opportunity cost of solving 1 Math question for Drake is: 10/15 = 2/3 Physics question.

b. The opportunity cost of solving 1 Physics question for Alice is 10/5 = 2 Math questions. The opportunity cost of solving 1 Physics question for Drake is 15/10 = 1.5 Math question.

c. Drake has the absolute advantage in doing both Physics homework and Math homework.

d. If they work together, then Alice will specialize in Math problems and Drake will specialize in Physics problems.

e. If they were to trade with each other, then the price range (in terms of number of Math problems) acceptable to both Alice and Drake for a Physics question is between 1.5 and 2 Math questions.